OpenStudy (anonymous):

A recipe for cupcakes uses 2 1/2 cups of sugar for 3 dozen cupcakes. How many cups of sugar must be used for 9 dozen cupcakes? Question 18 options: 4 5 1/2 7 1/2 27

OpenStudy (anonymous):

i think its A.

OpenStudy (anonymous):

It is definitely B because 9 is 3 times as large as 3 and 5 1/2 is 3 times as large as 2 1/2

OpenStudy (anonymous):

\[\frac{2.5 \text {cups}}{3 \text{ doz. cakes}} = \frac{\text{ ? cups }}{9\text{ doz. cakes }}\]

OpenStudy (anonymous):

Then solve for ?

OpenStudy (anonymous):

I don't get it .. im lost

OpenStudy (anonymous):

For three dozen, you need 2.5 cups of sugar.

OpenStudy (anonymous):

So \[\frac{2.5}{3}\] for cups over dozen.

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, then for 9 dozen you need ?? cups. where the ?? is what we want to know, but we can set up the fraction the same

OpenStudy (anonymous):

So, ?? cups over 9 dozen.

OpenStudy (anonymous):

so im basically adding 2.5 cups to 6?

OpenStudy (anonymous):

nope, we have two fractions, \[\frac{2.5}{3}\] and \[\frac{??}{9}\].

OpenStudy (anonymous):

So C 7 1/2 ?

OpenStudy (anonymous):

and is the ratio, or fraction of sugar to cupcakes stays the same, we can just set the two fraction equal.

OpenStudy (anonymous):

If we solve for ?, do we get 7.5? \[\frac{2.5 \text {cups}}{3 \text{ doz. cakes}} = \frac{\text{ ? cups }}{9\text{ doz. cakes }}\]

OpenStudy (anonymous):

Yea .thats what im getting

OpenStudy (anonymous):

\[2.5*9=?*3\]\[2.5*9/3=?\]\[2.5*3=?\]\[\implies ? = 7.5\]

OpenStudy (anonymous):

then, yes.

OpenStudy (anonymous):

2 more ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks